# Towards End-to-End Embodied Decision Making via Multi-modal Large Language Model: Explorations with GPT4-Vision and Beyond

### v1

@@ 1 Introduction
With the proposed benchmark, we conduct two series of evaluation: (1) We examine multiple state-of-the-art VLLMs, like InstructBLIP (Dai et al., 2023), MMICL (Zhao et al., 2023) and the latest GPT4-Vision (OpenAI, 2023a), in an end-to-end decision making context. (2) We introduce HOLMES, [^1] The system is aptly named after the renowned detective, Sherlock Holmes. a multi-agent cooperation framework. In this framework, we provide large language models, such as ChatGPT (OpenAI, 2022), GPT4 (OpenAI, 2023b), and Vicuna (Chiang et al., 2023), with descriptions of vision models like image captioning, object detection, Optical Character Recognition (OCR), and traffic sign detection models. Additionally, we supply descriptions of valid APIs within the simulated environment. The large language model subsequently initiates a search for clues pertaining to the question by engaging in a multi-turn conversation. This process involves alternating between invoking models or APIs to find clues and analyzing the discovered clues to facilitate informed decision making.
@@ 2 Related Work
Large language models pre-trained on large-scale multimodal (including text, image, video, etc.) corpus demonstrate impressive emergent capabilities and immense popularity (Brown et al., 2020; Wei et al., 2022), and have seen tremendous success in a wide range of domains covering various natural language processing and computer vision tasks (Radford et al., 2019; Chowdhery et al., 2022; Touvron et al., 2023; Alayrac et al., 2022; Zhu et al., 2023a; Li et al., 2023a). Consequently, using LLMs to empower the AI agents (Xi et al., 2023; Liu et al., 2023b; Park et al., 2023; Wang et al., 2023b; Yuan et al., 2023) becomes more and more promising. Specifically, we can employ LLMs to enhance decision-making ability of the agents (Nakano et al., 2022; Yao et al., 2022), expanding the perception and action space of these agents through strategies like tool utilization (Schick et al., 2023; Qin et al., 2023; Lu et al., 2023). Although LLM-based agents demonstrate reasoning and planning abilities through techniques like Chain of Thought or problem decomposition (Wei et al., 2023; Yao et al., 2023; Kojima et al., 2022), they inherently lack visual perception and are limited to only comprehending discrete textual content. Therefore, integrating visual information or other modalities can offer agents a broader context and a more precise understanding (Driess et al., 2023a), thereby enhancing their perception of the environment. However, no evaluation protocol or benchmark is currently available for evaluating decision making within the multimodal context.
@@ 3 PCA-EVAL
Unlike traditional visual question-answering datasets that emphasize visual perception (e.g., VQA (Goyal et al., 2017)), visual reasoning (e.g., NLVR (Suhr et al., 2017)), or world knowledge (e.g., OKVQA (Marino et al., 2019)), the most distinctive characteristic of PCA-EVAL is its grounding in embodied actions. Compared to embodied simulation environments such as ALFRED (Shridhar et al., 2020) and Minedojo (Fan et al., 2022), PCA-EVAL proves to be more effective in evaluating various LLM-based agents. This is primarily due to PCA-EVAL’s provision of high-level actions that can be readily implemented or programmed using the low-level actions within the corresponding domains. The high-level actions are more comprehensible for LLMs than the direct low-level actions like robotic movements in the simulation environments because, (1) the high-level actions are in the form of natural languages, which makes it easier for LLMs to understand the meaning and connect with world knowledge. (2) LLMs are not grounded with low-level actions during the pretraining or finetuning stage, making it hard for LLMs to understand the consequences of executing an action.
To answer a question in PCA-EVAL, the agent must possess the following abilities, (1) Perception: accurately identify the concept related to the question within the image, (2) Cognition: engage in reasoning based on image perception and worldly knowledge, and (3) Action: comprehend the potential actions, selecting the one that best aligns with the outcome of the reasoning process. A deficiency in any of these abilities would inevitably result in an incorrect answer, thereby posing a significant challenge to the more complex capabilities of embodied agents. Although challenging, all the aforementioned abilities are essential for the decision-making process in embodied environments.
@@ 3.1 Evaluation Metrics
The Action Score (A-Score) measures the model’s ability to generate appropriate and effective responses or actions based on the perceived input data and the cognitive understanding of the context. The score is assigned a value of 1 if the agent selects the correct action; otherwise, the score is set to 0. This score could be computed automatically.
The final Perception, Cognition, and Action scores of the agents are obtained by averaging the scores across all instances and domains in our PCA-EVAL dataset. Currently, perception and cognition scores are labeled by human annotators. The action score is computed automatically. We are working on making the full evaluation process human-free to unlock full features of the benchmark.
@@ 5.1 Configurations
We compare three different models, InstructBLIP-Vicuna-13B [^2] https://github.com/salesforce/LAVIS/tree/main/projects/instructblip, MMICL-FLANT5XXL [^3] https://huggingface.co/BleachNick/MMICL-Instructblip-T5-xxl and GPT4-Vision [^4] https://chat.openai.com. We apply the default inference configuration for the corresponding models.
@@
We compare three LLMs Vicuna [^5] https://huggingface.co/lmsys, ChatGPT-3.5-Turbo and GPT4666https://platform.openai.com. However we found Vicuna models lack the capability to call various APIs for information gathering, thus we have only reported the results for ChatGPT and GPT4. We anticipate supplementing these results as soon as open-source models become available, which can understand API descriptions and correspondingly call different APIs.
@@ 5.2 Evaluation
PCA-Eval assesses embodied decision-making through three distinct lenses: perception, cognition, and action. Currently, the action score is evaluated automatically using straightforward pattern-matching rules. However, the perception and cognition scores rely on the consensus score from three human evaluators. We compute the average kappa correlation coefficient for these evaluators, resulting in 0.66 for the Perception Score and 0.83 for the Cognition Score. These figures indicate a good consistency in the evaluation process. Looking ahead, our objective is to fully automate the evaluation of all scores, leveraging the powerful semantic understanding abilities of advanced LLMs. This advancement will enhance the effectiveness of PCA-Eval and simplify its usage.
@@ 5.3 Main Results
Table 1: Main results on PCA-EVAL. Models with $\dagger$ are fully open-source. Models with $\ddagger$ only provide API to access. P, C, and A represent Perception, Action, and Action Scores, respectively. For the open-source models in End2End setting, we find it hard to prompt them to output cross-modal reasoning information, so their Perception and Cognition scores are not reported.
@@
Figure 6: A Comparison between GPT4-V and GPT4-HOLMES
We also assessed the performance of embodied decision-making using our HOLMES system. As shown in the bottom block of the table, the HOLMES system, based on GPT4, achieves an Action Score of 0.71, matching the performance of GPT-4V (0.74). This suggests that the HOLMES system is proficient in understanding the task goal, breaking down the larger goal into multiple smaller steps, and accurately invoking the relevant APIs to accomplish each step.
@@ 6.1 Comparison Between End2End and HOLMES
Figure 7: An example of decision misalignment.
We conduct an analysis and comparison of the outputs generated by the End2End method with GPT4-Vision, as well as the HOLMES method with GPT4. Our findings indicate that the End2End method effectively mitigates information loss during the modality conversion process. As illustrated in Figure 6, an image depicts a road with several nearby cars. GPT4-Vision is capable of discerning that these cars are situated in a safe space, thereby suggesting that the driver can continue driving.
@@ 6.2 Alignment between Agent Decisions and Human Values
We have observed instances where the decisions made by the agent contradict human values. For instance, consider the scenario depicted in Figure 7. The image illustrates a crosswalk devoid of pedestrians. The appropriate response in this situation would be to slow down, as caution is paramount when approaching a crosswalk, regardless of the presence or absence of pedestrians. However, upon processing the information that the crosswalk is unoccupied, ChatGPT suggests that maintaining the current speed is the optimal action, arguing that the absence of pedestrians eliminates the need to slow down. The rationale provided by ChatGPT is logical, yet it does not align with human values. We believe it is crucial for embodied agents to make decisions that are in harmony with human values, rather than solely focusing on maximizing their advantage.
@@ 6.3 Limitation and Future Work
At present, our evaluation baseline is capable of automatically assessing only the action score. Moving forward, we aspire to develop automated evaluation systems for both perception and cognition scores. This advancement will significantly augment the ease and efficiency of future evaluations.

### v4

@@ 1 Introduction
With the proposed benchmark, we conduct two series of evaluation: (1) We examine multiple state-of-the-art VLLMs, like InstructBLIP (Dai et al., 2023), MMICL (Zhao et al., 2023), QwenVL-Chat (Bai et al., 2023) and the latest GPT4-Vision (OpenAI, 2023a), in an end-to-end decision making context. (2) We introduce HOLMES, [^1] The system is aptly named after the renowned detective, Sherlock Holmes. a multi-agent cooperation framework. In this framework, we provide large language models, such as ChatGPT (OpenAI, 2022), GPT4 (OpenAI, 2023b), and Vicuna (Chiang et al., 2023), with descriptions of vision models like image captioning, object detection, Optical Character Recognition (OCR), and traffic sign detection models. Additionally, we supply descriptions of valid APIs within the simulated environment. The large language model subsequently initiates a search for clues pertaining to the question by engaging in a multi-turn conversation. This process involves alternating between invoking models or APIs to find clues and analyzing the discovered clues to facilitate informed decision making.
@@ 2 Related Work
Large language models pre-trained on large-scale multimodal (including text, image, video, etc.) corpus demonstrate impressive emergent abilities and immense popularity (Brown et al., 2020; Wei et al., 2022), and have seen tremendous success across various domains covering various natural language processing and computer vision tasks (Radford et al., 2019; Chowdhery et al., 2022; Touvron et al., 2023; Alayrac et al., 2022; Zhu et al., 2023a; Li et al., 2023a). Consequently, using LLMs to empower the AI agents (Xi et al., 2023; Liu et al., 2023b; Park et al., 2023; Wang et al., 2023d; Yuan et al., 2023) becomes more and more promising. Specifically, we can employ LLMs to enhance the decision making ability of the agents (Nakano et al., 2022; Yao et al., 2022; Li et al., 2023c; Song et al., 2023), expanding their perception and action space through strategies like tool utilization (Schick et al., 2023; Qin et al., 2023; Lu et al., 2023). Although LLM-based agents demonstrate reasoning and planning abilities through techniques like Chain of Thought or problem decomposition (Wei et al., 2023; Yao et al., 2023; Kojima et al., 2022), they inherently lack visual perception, and are limited to the discrete textual content. Therefore, integrating visual information or other modalities can offer agents a broader context and a more precise understanding (Driess et al., 2023a), enhancing their environmental perception. However, no evaluation protocol or benchmark is currently available to evaluate decision making within the multimodal context.
@@ 3 PCA-EVAL
Unlike traditional visual question-answering datasets that emphasize visual perception (e.g., VQA (Goyal et al., 2017)), visual reasoning (e.g., NLVR (Suhr et al., 2017)), or world knowledge (e.g., OKVQA (Marino et al., 2019)), the most distinctive characteristic of PCA-EVAL is its grounding in embodied actions. Compared to embodied simulation environments like ALFRED (Shridhar et al., 2020) and Minedojo (Fan et al., 2022), PCA-EVAL proves to be more effective in evaluating various LLM-based agents. This is primarily due to PCA-EVAL’s provision of high-level actions that can be readily implemented or programmed using the low-level actions in the corresponding domains. The high-level actions are more comprehensible for LLMs than the direct low-level actions like robotic movements in the simulation environments because (1) the high-level actions are in the form of natural languages, making it easier for LLMs to understand the meaning and connect with world knowledge. (2) LLMs are not grounded with low-level actions during the pretraining or finetuning stage, making it hard for LLMs to understand the consequences of executing an action.
To answer a question in PCA-EVAL, the agent must possess the following abilities: (1) Perception: accurately identify the concept related to the question within the image; (2) Cognition: engage in reasoning based on image perception and worldly knowledge; (3) Action: comprehend the potential actions, selecting the one that best aligns with the outcome of the reasoning process. A deficiency in any of these abilities would inevitably result in an incorrect answer, posing a significant challenge to the more complex capabilities of embodied agents. Although challenging, all the aforementioned abilities are essential for the decision-making process in embodied environments.
@@ 3.1 Evaluation Metrics
The Action Score (A-Score) measures the model’s ability to generate appropriate and effective responses or actions based on the perceived input data and the cognitive understanding of the context. The score is assigned a value of 1 if the agent selects the correct action; otherwise, the score is set to 0.
The final Perception, Cognition, and Action scores of the agents are obtained by averaging the scores across all instances and domains in our PCA-EVAL dataset.
3.2 Automatic Evaluation
Recent advancements have seen researchers harnessing powerful LLMs for the evaluation of output of language models. Studies have revealed that the outcomes from LLMs could exhibit remarkable alignment with human judgments Zheng et al. (2023); Wang et al. (2023c; b). In our investigation, we employed GPT-4 to automatically evaluate perception, cognition, and action scores based on the model’s outputs. Our findings underscore a significant agreement between GPT-4 annotations and human annotator results. This is substantiated by Pearson correlation coefficients of 0.8, 0.9, and 0.95 for perception, cognition, and action evaluations, respectively. To facilitate ongoing and future research endeavors, we share our automatic evaluation script [^2] https://github.com/pkunlp-icler/PCA-EVAL/blob/main/pca-eval/evaluation/pca_auto_scoring.py for seamless adoption, which could also be improved in the future. For a detailed description of our evaluation methodology, kindly refer to Appendix C
@@ 5.1 Configurations
We compare four different models, InstructBLIP-Vicuna-13B [^3] https://github.com/salesforce/LAVIS/tree/main/projects/instructblip, MMICL-FLANT5XXL [^4] https://huggingface.co/BleachNick/MMICL-Instructblip-T5-xxl, QwenVL-Chat [^5] https://huggingface.co/Qwen/Qwen-VL-Chat and GPT4-Vision [^6] https://chat.openai.com. We apply default inference configurations for the corresponding models.
@@
We compare three LLMs: Vicuna [^7] https://huggingface.co/lmsys, ChatGPT-3.5-Turbo and GPT4888https://platform.openai.com. However we found Vicuna models lack the capability to call various APIs for information gathering, thus we have only reported the results for ChatGPT and GPT4. We anticipate supplementing these results as soon as open-source models become available, which can understand API descriptions and correspondingly call different APIs.
@@ 5.2 Evaluation
PCA-Eval assesses embodied decision-making through three distinct lenses: perception, cognition, and action. The scores we reported in Table 1 rely on the consensus score from three human evaluators. We compute the average kappa correlation coefficient for these evaluators, resulting in 0.91 for the Perception Score and 0.88 for the Cognition Score. These figures indicate a good consistency in the evaluation process.
@@ 5.3 Main Results
Table 1: Main results on PCA-EVAL. Models with $\dagger$ are fully open-source. Models with $\ddagger$ only provide API to access. P, C, and A represent Perception, Cognition, and Action Scores, respectively. For the open-source models in End2End setting, we find it hard to prompt them to output correct cross-modal reasoning information, so their Perception and Cognition scores are not reported.
@@
We also assessed the performance of embodied decision making using our HOLMES system.
As shown in the bottom block of the table, the HOLMES system, based on GPT4, achieves an Action Score of 0.71, matching the performance of GPT-4V (0.74). This suggests that the HOLMES system is proficient in understanding the task goal, breaking down the larger goal into multiple smaller steps, and accurately invoking the relevant APIs to accomplish each step.
@@ 6.1 Comparison Between End2End and HOLMES
We conduct an analysis and comparison of the outputs generated by the End2End method with GPT4-Vision, as well as the HOLMES method with GPT4. Our findings indicate that the End2End method effectively mitigates information loss during the modality conversion process. As illustrated in Figure 6a, an image depicts a road with several nearby cars. GPT4-Vision is capable of discerning that these cars are situated in a safe space, thereby suggesting that the driver can continue driving.
@@
(a) A Comparison between GPT4-V and GPT4-HOLMES
(b) An example of decision misalignment.
Figure 6: Case studies.
@@ 6.2 Alignment between Agent Decisions and Human Values
We have observed instances where the decisions made by the agent contradict human values. For instance, consider the scenario depicted in Figure 6b. The image illustrates a crosswalk devoid of pedestrians. The appropriate response in this situation would be to slow down, as caution is paramount when approaching a crosswalk, regardless of the presence or absence of pedestrians. However, upon processing the information that the crosswalk is unoccupied, ChatGPT suggests that maintaining the current speed is the optimal action, arguing that the absence of pedestrians eliminates the need to slow down. The rationale provided by ChatGPT is logical, yet it does not align with human values. We believe it is crucial for embodied agents to make decisions that are in harmony with human values, rather than solely focusing on maximizing their advantage.
@@ 11    Lists items in your hand, which you don’t need to pick up
Appendix C Automatic Evaluation
[Question]: {question} [Action Choices]: {actions} [Agent Answer]: {model_output} [Correct Action]: {true_action} [Key Concepts]: {key_concept} [Reference Reasoning Process]: {reason} [System] We would like you to access the agent’s performance in the multimodal reasoning task about domain. In this task, the agent is given an image, a [Question], and several candidate [Action Choices], and is asked to give an [Agent Answer] for the [Question]. The [Agent Answer] encapsulates the agent’s perception of the image’s [Key Concepts], the agent’s cognition reasoning process and the final selected action. We request you to give three types of scores for the agent’s [Agent Answer] in comparison to the given [Key Concepts], [Reference Reasoning Process] and [Correct Action]: 1. action score: If the selected action in the [Agent Answer] matches that of the [Correct Action], the action score is 1; otherwise, it is 0. 2. perception score: This score evaluates the model’s capability to perceive and interpret observations. It is contingent on whether the [Agent Answer] includes any of the [Key Concepts] of the instance. If it accurately describes any one of the [Key Concepts], the score is 1; otherwise, it is 0. 3. cognition score: This score gauges the model’s ability to reason, comprehend, and make informed decisions based on perceived input data and world knowledge. If the reasoning process in the [Agent Answer] aligns with the [Reference Reasoning Process], the score is 1; otherwise, it is 0. Please note that there are only scores of 0 and 1. You should carefully compare the [Agent Answer] with the [Correct Action], [Key Concepts] and [Reference Reasoning Process] to give your assessment. You need first to give your assessment evidence and then the scores. Your output MUST contain 6 lines with the following format: action assessment evidence: (assessment evidence here) action score: (score here) perception assessment evidence: (assessment evidence here) perception score: (score here) cognition assessment evidence: (assessment evidence here) cognition score: (score here)
Table 3: We utilize the template to query GPT-4, aiming to evaluate its responses and assign scores for perception, cognition, and action. By feeding both the agent’s output and the ground truth answer to GPT-4, based on this template, we can then extract the three distinct scores from the conclusion of GPT-4’s response.